Convert 14.0 moles of methane, CH4, to molecules.


1
Expert's answer
2021-03-02T02:01:13-0500

N(CH4) = n(CH4)*NA = (14.0 mol)*(6.02*1023 mol-1) = 8.428*1024 molecules